Wendy and Charlotte both enjoy watching Manu’s Kitchen Rules (or MKR). Initially, assume that when broadcast on public television MKR is a public good that is both non-excludable and non-rivalrous. Wendy and Charlotte have demand for MKR given by the following:

Qw=122 − 2

0 0
Answer #1

Ansnden a) Market demand Q I6 2-3P Maxgival Cost unden Compe ttfve econamy. P -mc = 50 Quavitity puce d q= 162-3 (50) = 11 Eb